OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. B(0, 0), C(4, -4)

OpenStudy (anonymous):



OpenStudy (anonymous):

All right, here is our graph with points B and C shown. Now, can you show me an equation for a line? \[y=m*x + b\] is common equation for a line.

OpenStudy (anonymous):

You can determine the slope, m. Then you can determine the y-intercept which is what b is.

OpenStudy (anonymous):

You can compute the slope by \[ m= \frac{y2 - y1}{x2 - x1}\] Where you have two points: B and C B = (0,0) = (x1, y1) C = (4, -4) = (x2, y2) Therefore, using these values to obtain the slope, m: \[ m= \frac{(-4) - 0}{4 - 0}\] Now simplify that.

OpenStudy (anonymous):

Once you have m, plug it in the equation \[y=m*x + b\]. Then you can use one of the points to determine what b is in the equation: Here, using point B, where B = (0,0) = (y,x) 0 = m*0 + b Solve for b

OpenStudy (anonymous):

That is an approach that should get you the answer: give it a try, and I can check your work for you once you have a solution ;)

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

No problem, I hope it is clear. Please ask if you have any questions or need clarification.

OpenStudy (anonymous):

ok sounds good thanks again :)